# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ E.4 Darcy flow forward operator learning
That is, $\mathcal{G}^{\star}\colon a\mapsto u$. Here, $\Omega\coloneqq(0,1)^{2}$.
Again, we first assign a random measure $\mathbb{Q}$ to evaluate the OOD performance. We take $\mathbb{Q}=\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu_{k}^{\prime}}$, where each test distribution $\nu_{k}^{\prime}$ is defined by parameters sampled i.i.d. from custom distributions (E.13): $m_{k}\sim\rho([0,1])$, $\alpha_{k}\sim\rho([2,3])$, and $\tau_{k}\sim\rho([1,4])$. The scaling parameter is set as $\sigma_{k}=\tau_{k}^{\alpha_{k}-1}$. We set $K=20$ and the grid parameter $M=64$. Given these parameters, each random conductivity function $a^{\prime}$ from test distribution $\nu_{k}^{\prime}$ has the form
$x\mapsto a^{\prime}(x)=\exp\biggl{(}m_{k}+\sum_{i=1}^{M+1}\sum_{j=1}^{M+1}\frac{2\sigma_{k}}{\bigl{(}(i\pi)^{2}+(j\pi)^{2}+\tau_{k}^{2}\bigr{)}^{\alpha_{k}/2}}z^{\prime}_{ij}\sin(i\pi x_{1})\sin(j\pi x_{2})\biggr{)}\,,$ (E.22)
where $z^{\prime}_{ij}\sim\mathcal{N}(0,1)$ independently. Sample conductivities $a$ from training distribution $\nu$ are assumed to follow the same form
$x\mapsto a(x)=\exp\biggl{(}m+\sum_{i=1}^{M+1}\sum_{j=1}^{M+1}\frac{2\sigma}{\bigl{(}(i\pi)^{2}+(j\pi)^{2}+\tau^{2}\bigr{)}^{\alpha/2}}z_{ij}\sin(i\pi x_{1})\sin(j\pi x_{2})\biggr{)}$ (E.23)
with fixed parameters $\alpha=2$, $\tau=3$, and $\sigma=3$ and $z_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. Alg. 2 is used to optimize the mean shift parameter $m\in\mathbb{R}$ in order to reduce the OOD generalization error. As in SM E.3, $\log(a)$ is modeled via a finite KLE of a Gaussian measure with Matérn-like covariance operator. Thus, by working with $\log(a)$ directly, we can exploit (A.1) to write the squared $2$-Wasserstein distance factor in objective (4.7) in closed-form.
@@
The same DeepONet architecture and training setup from SM E.3 are used here, except the branch network has layer sizes $M^{2}$, $256$, $256$, $128$, and $200$. After the model-training half step in Alg. 2, the scalar parameter $m$ is optimized using PyTorch’s Adam optimizer with a learning rate of $10^{-3}$.
@@
The initial value of $m$ is set to zero. The training and optimization procedure follows the same structure as in SM E.3, but with the following differences: $200$ conductivity–solution pairs are generated for training using the current value of $m$. The model is initially trained for $5000$ iterations. If the AMA loss has not improved relative to the previous iteration, training continues for an additional $5000$ iterations. This process is repeated up to three times, after which the algorithm proceeds to the distribution parameter optimization step. During distribution parameter optimization, $500$ steps are taken. Then the AMA loss is evaluated using three newly generated data pairs. If the AMA loss has not improved relative to the previous iteration, another $500$ steps are taken, up to $1500$ total steps. After this, the procedure returns to model training again. To ensure deterministic behavior, $203$ sets of random coefficients $\{z_{ij}\}$ are fixed across training and optimization. The preceding process is repeated across $13$ independent runs, each using different realizations of the coefficients. The average relative OOD error and average relative AMA loss are computed for each run.
@@
After $13$ independent runs, the average relative OOD error and relative AMA loss are computed and are shown in Fig. 8. In Fig. 4, we visualize the pointwise absolute error for iterations $1$, $2$, $3$, and $8$. At these iterations, the DeepONet model is trained using the optimal training distribution produced by AMA. The model is then evaluated on the same fixed test conductivity. The progressive reduction in the spatial error across these panels of Fig. 4 corroborates the quantitative decrease in the relative OOD error, which converges to approximately $4\%$ in our final model.
Figure 8: Average relative OOD error and average relative loss of Alg. 2 on the Darcy flow forward problem. (Left) After $13$ independent runs, decay of the average relative OOD error of the model when trained on the optimal distribution identified at each iteration; a $95$% confidence interval of the true relative OOD error is provided at each iteration. (Right) Decay of average AMA loss defined in (4.7) vs. iteration relative to the same loss at initialization.
@@
The computational resources here are the same as those reported in SM E.3, except three total CPU and GPU hours were needed for all independent runs.

### v3

@@ F.5 Radiative transport operator learning
where $a$ is the scattering coefficient, $\varepsilon$ is the Knudsen number, and the collision operator $u\mapsto Q[u]$ is given by
$Q[u](z,v)\coloneqq\int_{V}\mathcal{K}(v,v^{\prime})u(z,v^{\prime})\,dv^{\prime}-u(z,v)\,.$ (F.25)
In this example, we use the constant kernel $\mathcal{K}(v,v^{\prime})=\tfrac{1}{|V|}$. The inflow boundary is defined as
$\Gamma_{-}\coloneqq{\mathopen{}\mathclose{{\left\{(z,v)\in\partial\Omega\times V\,\middle|\,-\bm{n}_{z}\cdot v\geq 0}}\right\}}\,$, (F.26)
where $\bm{n}_{z}$ denotes the outward normal at $z\in\partial\Omega$.
On three separate experiments with error results given in Fig. 5, we set the Knudsen number $\varepsilon$ equal to 1/8, 2, and 8. Changing $\varepsilon$ puts us in different physical regimes as seen in Fig. 17. The inflow boundary condition is given by
$\phi(z,v)=\begin{cases}\bigl(1+0.5\sin(\pi y/L)\bigr)\max(v\cdot\bm{n}_{z},0),&\text{if }\ z=0\,,\\ 0,&\text{elsewhere on }\Gamma_{-}\,.\end{cases}$ (F.27)
Figure 17: Mapping from scattering coefficient to spatial density in the radiative transport equation for three Knudsen numbers $\varepsilon$. Larger $\varepsilon$ values yield a more hyperbolic regime.
We seek the optimal input training distribution $\nu$ for learning the nonlinear operator that maps the scattering coefficient function $a(\cdot)$ to the spatial domain density $\int_{V}u(\cdot,v)\,dv$.
In this example, $\Omega\coloneqq(0,1)^{2}$ is the unit square and $V\coloneqq\mathbb{S}^{1}={\mathopen{}\mathclose{{\left\{(\cos\theta,\sin\theta)\,\middle|\,\theta\in[-\pi,\pi)}}\right\}}$ is the unit circle. Thus, we parametrize $v$ by $\theta$, i.e., $v(\theta)=(\cos\theta,\sin\theta)$ for $\theta\in[-\pi,\pi)$, $\int_{V}u(\cdot,v)\,dv=\int_{-\pi}^{\pi}u(\cdot,v(\theta))\,d\theta$, and $v\cdot\nabla_{z}=(\cos\theta)\partial_{x_{1}}+(\sin\theta)\partial_{x_{2}}$. We first generate samples from $\mathbb{Q}$ to then evaluate the OOD performance of the model trained on $\nu$. We set $K=3$. Each distribution $\nu_{k}^{\prime}\sim\mathbb{Q}$ is set to be $(\exp)_{\#}\gamma_{k}$, where $\gamma_{k}$ is a Gaussian process parametrized by $\sigma_{k}$, $\alpha$, and $\tau$. Thus, samples from $\nu_{k}^{\prime}$ have the form
$a^{\prime}(x)=\exp\biggl(f_{k}(x)+\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{2\sigma_{k}}{\bigl((i\pi)^{2}+(j\pi)^{2}+\tau^{2}\bigr)^{\alpha/2}}z_{ij}\sin(i\pi x_{1})\sin(j\pi x_{2})\biggr)\,$, (F.28)
where $z_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. The mean $f_{k}$ of each distribution $\gamma_{k}$ is selected as follows. For a fixed centered Gaussian process $\gamma$, draw $f_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\gamma$ for each $k$. Then these $f_{k}$ functions are then fixed once and for all at the beginning of the experiment. By supplying distinct nonzero means to each of the $K$ test distributions, the problem of finding the optimal training distribution becomes harder. Indeed, the test distributions now exhibit not only covariance shifts, but also mean shifts. We set $\tau=1$ and $\alpha=2$ for all $K=3$ test distributions, and set $\sigma_{1}=1$, $\sigma_{2}=2$, and $\sigma_{3}=3$. Note that the parameters of $\gamma$ are set to $\tau=1$, $\alpha=2$, and $\sigma=1$.
@@
Following Alg. 2, we represent $\nu$ as an empirical measure of $N$ equally weighted particles: $\nu=\frac{1}{N}\sum_{q=1}^{N}\delta_{a_{q}}$. Each particle $a_{q}$ is a function of the form
$a_{q}(x)=\exp\mathopen{}\mathclose{{\left(\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}c_{qij}\cdot 2\sin(i\pi x_{1})\sin(j\pi x_{2})}}\right)\,$, (F.29)
where the coefficients $c_{q}:=\{c_{qij}\}_{i,j\geq 1}$ belong to $\ell^{2}$. Rather than optimizing the particles $a_{q}$ directly, we optimize their representations in the coefficient space through the transformation
$a_{q}=\exp(Lc_{q})\,$, (F.30)
where $L\colon\ell^{2}\to L^{2}(\Omega)$ is the linear operator mapping coefficients to the series in (F.29) and $\exp$ is applied pointwise. The distribution $\nu$ is thus obtained as the pushforward $\nu=(\exp\circ L)_{\#}\mu$, where $\mu$ is the distribution over coefficient sequences. We can then write the loss as $\widetilde{F}(\mu)\coloneqq F((\exp\circ L)_{\#}\mu)=F(\nu)$, where $F$ is defined as the upper bound in Prop. 3.1, i.e., $\widetilde{F}=F\circ(\exp\circ L)_{\#}$. For each step $k$, the discretized Wasserstein gradient flow update rule with respect to $\{c_{q}\}_{q=1}^{N}$ is then
$c_{q}^{(k+1)}=c_{q}^{(k)}-\eta^{(k)}\nabla_{\ell^{2}}\Bigl[D\widetilde{F}\bigl(\mu^{(k)}_{N}\bigr)\Bigr]\bigl(c_{q}^{(k)}\bigr)\,,\quad\text{where}\quad\mu^{(k)}_{N}\coloneqq\frac{1}{N}\sum_{q^{\prime}=1}^{N}\delta_{c_{q^{\prime}}^{(k)}}$ (F.31)
and $\eta^{(k)}$ is the step size. By the chain rule and linearity of pushforward,
$D\widetilde{F}(\mu)=DF\bigl((\exp\circ L)_{\#}\mu\bigr)\circ(\exp\circ L)$ (F.32)
for each $\mu$, where $DF$ denotes the derivative from Prop. 4.9. Thus, we can write the update rule (F.31) in terms of the known derivative $DF$.
The $\ell^{2}$ gradient with respect to $c_{q}$ is computed by a combination of PyTorch’s autograd and the adjoint state method. After updating $\{c_{q}\}_{q=1}^{N}$, we train a model with discretized versions of particles $\{a_{q}\}_{q=1}^{N}$ denoted as $\{\hat{a}_{q}\}_{q=1}^{N}$. Each particle $a_{q}$ is discretized as an $m\times m$ matrix. Similar to the setups from SM F.3 and SM F.4, the model architecture is a DeepONet. However, the branch network now has $6$ fully connected layers of sizes $m^{2}$, $100$, $100$, $100$, $64$, and $32$, and the trunk network has $6$ fully connected layers of sizes $2$, $100$, $100$, $100$, $64$ and $32$.
@@
During optimization, we work with $N$ samples from each distribution $\{\nu^{\prime}_{k}\}_{k=1}^{K}$ rather than the distributions themselves. For computational simplicity, we set the number of samples from each $\nu^{\prime}_{k}$ equal to the number of particles in the training distribution.
The particles $\{a_{q}\}_{q=1}^{N}$ are discretized by truncating the number of terms in its expansion, i.e.,
$a_{q}(x)=\exp\mathopen{}\mathclose{{\left(\sum_{i=1}^{N_{\text{truncate}}}\sum_{j=1}^{N_{\text{truncate}}}c_{qij}\cdot 2\sin(i\pi x_{1})\sin(j\pi x_{2})}}\right)\,.$ (F.33)
The coefficient sequences $\{c_{q}\}_{q=1}^{N}$ are truncated to $N_{\text{truncate}}=20$ for both indices $i$ and $j$, yielding sequences of 100 coefficients each. These are initialized by sampling from $\mathcal{N}(0,10^{-6})$.
We discretize $\Omega$ on a $51\times 51$ grid ($m=51$) and $V$ with $12$ different directions, i.e. $\Delta\theta=\pi/6$, so we approximate the collision term by
$Q[u_{q}]$ $=\frac{1}{|V|}\int_{V}u_{q}(z,v^{\prime})\,dv^{\prime}-u_{q}(z,v)$
$\approx\frac{1}{|V|}\sum_{i=1}^{12}u_{q}(\cdot,\theta_{i})\,\Delta\theta-u_{q}$
$=\frac{\pi/6}{2\pi}\sum_{i=1}^{12}u_{q}(\cdot,\theta_{i})-u_{q}$
$=\frac{1}{12}\sum_{i=1}^{12}u_{q}(\cdot,\theta_{i})-u_{q}$
and the spatial-domain density by
$\tilde{u}_{q}\coloneqq\int_{V}u_{q}(\cdot,v)\,dv\approx\sum_{i=1}^{12}u_{q}(\cdot,\theta_{i})\,\Delta\theta\,.$ (F.34)
Initialize the step size as $\eta=1\times 10^{-6}$. At each iteration of Alg. 2:
  1. 1.
    Compute the true outputs $\{\tilde{u}_{q}\}_{q=1}^{N}$ as $m\times m$ matrices using the current $\{c_{q}\}_{q=1}^{N}$.
  2. 2.
    Train the DeepONet for 5,000 epochs with an 80/20 train-test split.
  3. 3.
    Update $\{c_{q}\}_{q=1}^{N}$ according to Eqn. (F.31).
  4. 4.
    If the loss $F(\nu)$ increases, halve the step size and recompute the update. The decreased step size carries over to the next iteration.
[…]
For the loss computation in Step 4 above, the second moment of a distribution $\nu$ is estimated with $N$ samples by
$\mathsf{m}_{2}(\nu)=\operatorname{\mathbb{E}}_{a\sim\nu}\|a\|_{L^{2}(\Omega)}^{2}\approx\frac{1}{N}\sum_{i=1}^{N}\|a_{i}\|_{L^{2}(\Omega)}^{2}\,$, (F.35)
where $\|a_{i}\|_{L^{2}(\Omega)}=\int_{\Omega}a_{i}(z)^{2}dz$ is estimated with the trapezoidal rule. Note that for our empirical training distribution, F.35 is not an approximation, but instead an equality. We estimate the Wasserstein distance $\mathsf{W}_{2}(\nu,\nu^{\prime})$ between distributions $\nu$ and $\nu^{\prime}$ using $N$ samples from each, with the ground cost function taken to be $c(a,a^{\prime})=\|a-a^{\prime}\|^{2}_{L^{2}(\Omega)}$.
@@
We compare models trained on the optimized distribution at each iteration of particle-based AMA with a benchmark model trained on $N$ samples from the mixture of test distributions $\nu_{\mathbb{Q}}=\frac{1}{K}\sum_{k=1}^{K}\nu_{k}^{\prime}$ (Eqn. D.1), using the same DeepONet architecture and 5,000 training epochs for all models. Performance is measured by the relative OOD error (Eqn. F.19), approximated with the original $N$ samples from each $\nu_{k}^{\prime}$. As shown in Fig. 5, the optimized model improves substantially after particle-based AMA, achieving performance comparable to the benchmark and even outperforming the benchmark in small sample-size regimes. Given that AMA minimizes an upper bound, we are not directly minimizing the OOD error which can account for why the optimized distribution does not outperform the benchmark distribution for most sample sizes. Tab. 3 reports 95% confidence intervals for relative OOD errors over 10 independent trials at $N=120$, demonstrating that the optimized model reduces its relative OOD error by 82% from its original value on average for $\varepsilon=\tfrac{1}{8}$, 85% for $\varepsilon=2$, and 88% for $\varepsilon=8$. For this particular sample size, the runtime, shown in Fig. 18, averages 20.6 minutes for $\varepsilon=\tfrac{1}{8}$, 26.7 minutes for $\varepsilon=2$, and 26.1 minutes for $\varepsilon=8$.
Table 3: Radiative transport operator learning with particle-based Alg. 2. Reported are mean relative OOD errors with 95% confidence intervals computed over 10 independent runs for three separate models with varying Knudsen number $\varepsilon$. All models have the same DeepONet architecture and are trained with $N=120$ samples for 5,000 epochs. What makes each model different is their training distribution. The initial model was trained with random particles $c_{qij}\sim\mathcal{N}(0,10^{-6})$, the optimized model was trained with optimized particles from particle-based AMA, and the benchmark model is a model trained with samples from the mixture of test distributions, i.e., $\nu_{\mathbb{Q}}=\frac{1}{K}\sum_{k=1}^{K}\nu_{k}^{\prime}$.
$\varepsilon$ Initial Optimized Benchmark
$1/8$ $8.65\times 10^{-2}\pm 8.18\times 10^{-3}$ $1.55\times 10^{-2}\pm 9.72\times 10^{-4}$ $1.62\times 10^{-2}\pm 5.34\times 10^{-3}$
22 $8.76\times 10^{-2}\pm 8.63\times 10^{-3}$ $1.29\times 10^{-2}\pm 1.20\times 10^{-3}$ $1.24\times 10^{-2}\pm 2.69\times 10^{-3}$
88 $9.09\times 10^{-2}\pm 4.39\times 10^{-3}$ $13\times 10^{-2}\pm 1.33\times 10^{-3} 1.1$$16\times 10^{-2}\pm 2.26\times 10^{-3}$
Figure 18: Runtime of particle-based AMA for the radiative transport equation with sample size NN at three Knudsen numbers $\varepsilon\in\{1/8,2,8\}$. Error bars indicate 95% confidence intervals over 10 trials.
@@
All experiments were conducted on AWS using Galaxy’s Deep Learning AMI GPU, PyTorch 2.0.1, and Ubuntu 20.04-prod-ada4rwvf26mda with an EC2 instance type of g5.12xlarge. The experiments for the three different $\varepsilon$ values were run in parallel, with the $\varepsilon=2$ case taking the longest at 27 hours. Thus, a total of approximately 27 GPU-hours (and corresponding CPU usage) were required for all independent runs with three GPUs running in parallel.
F.6 Burgers’ equation operator learning
In the final example from Sec. 5, we consider the viscous Burgers’ equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\kappa\frac{\partial^{2}u}{\partial x^{2}}$ (F.36)
with viscosity $\kappa=0.1$ on the periodic domain $x\in[0,2\pi]_{\mathrm{per}}$. Our goal is to learn the operator that maps the initial condition $a\coloneqq u({\,\cdot\,},0)$ to the solution $u\coloneqq u({\,\cdot\,},1)$. This benchmark problem is taken from Nelsen and Stuart [72], Li et al. [57], Kovachki et al. [52].
Following Alg. 2, we model the training data distribution $\nu$ as a Gaussian process with a Matérn covariance kernel. Samples from $\nu$ are generated according to
$a(x)=\frac{1}{\sqrt{\pi}}\sum_{j=1}^{n/2}\frac{\sigma}{(j^{2}+\tau^{2})^{\alpha/2}}\bigl(z^{(s)}_{j}\sin(jx)+z^{(c)}_{j}\cos(jx)\bigr)\,$, (F.37)
where $z^{(s)}_{j}$ and $z^{(c)}_{j}$ are i.i.d. samples from $\mathcal{N}(0,1)$ and $n=8192$ is the discretization size. Only $n/2$ modes are retained due to the Nyquist limit.
Model architecture and optimization algorithm.
We adopt the DeepONet architecture described in SM F.3. The branch network has layer sizes $n$, $1024$, $512$, $256$, $128$, $p$ and the trunk network has layer sizes $1$, $128$, $128$, $128$, $128$, $p$ where $p=128$. Training uses Adam with initial learning rate $10^{-3}$ and inverse-time decay (decay steps = 2000,  decay rate = 0.9). We add $\ell_{2}$ regularization via weight decay $10^{-3}$ and apply early stopping on the test loss with patience 1000 and $\mathrm{min\_delta}=10^{-4}$. The maximum number of training iterations is 10,000. After the model-training half-step in Alg. 2, we optimize the Matérn covariance parameters $(\alpha,\tau,\sigma)$ by minimizing the upper bound in Prop. 3.1 using scipy.optimize.minimize, with gradients from Lem. D.8.
Training and optimization procedure.
We initialize the parameters at $\alpha=3$, $\tau=1.4$, and $\sigma=5$. All experiments use the fixed publicly available dataset burgers_data_R10.mat from Kovachki et al. [52], which contains 2048 input-output sample pairs. We reserve $10\%$ of the data (205 samples) as a held-out set for estimating the relative OOD error and treat the remaining 1843 samples as i.i.d. draws from a test distribution which we assume to be Gaussian. To optimize the training distribution $\nu$, we generate $N=1843$ samples from the current parameter values, train the DeepONet, and then update the parameters using this model and test samples. We repeat this process until the improvement in the objective function falls below $3\times 10^{-3}$. This occurs at iteration $10$, yielding optimal parameters $\alpha=2.906$, $\tau=1.456$, $\sigma=5.005$.
Model evaluation.
After identifying the optimal training distribution, we compare the model trained on $N$ samples from this distribution with models trained using two pool-based active learning strategies, Query-by-Committee (QbC) and CoreSet. We treat the 1843 test samples as the pool, i.e., $N_{\mathrm{pool}}=1843$. Each method begins with 10 randomly selected samples, after which we fully train the model, select an additional $n_{\mathrm{select}}=40$ samples according to the method-specific criterion, and retrain the model from scratch. For subsequent iterations we use $n_{\mathrm{select}}=50$ until each method reaches a total of 1000 training samples. We run 10 independent trials of this experiment. The selection criteria for each strategy are summarized below.
Selection criteria.
The pool contains $N_{\mathrm{pool}}$ candidate inputs. Let $n_{\mathrm{select}}$ denote the number of points selected at each acquisition step and let $\Delta x=2\pi/n$ be the spatial grid spacing.
  • Query-by-Committee (QbC). Given an ensemble of $M=4$ models that produce predictions $u_{i}^{(m)}\in\mathbb{R}^{n}$ for pool sample $i$ and model $m$, define the ensemble mean $\bar{u}_{i}=\tfrac{1}{M}\sum_{m=1}^{M}u_{i}^{(m)}$. The uncertainty score is the average squared $L^{2}$ deviation from the mean:
    $\mathrm{uncertainty}(i)\coloneqq\frac{1}{M}\sum_{m=1}^{M}\big\|u_{i}^{(m)}-\bar{u}_{i}\big\|_{L^{2}}^{2}\quad\text{and}\quad\big\|v\big\|_{L^{2}}\coloneqq\sqrt{\Delta x\sum_{k=1}^{n}v_{k}^{2}}\,.$ (F.38)
    Select the $n_{\mathrm{select}}$ pool points with largest $\mathrm{uncertainty}(i)$.
  • CoreSet (greedy $k$-center on last-layer features). Let $k:=n_{\mathrm{select}}$. For each pool input, compute a last-layer feature vector $f_{i}\in\mathbb{R}^{p}$ from the DeepONet evaluated at point $a_{i}$ from the pool, apply a Gaussian sketch with $d=32$, i.e.,
    $U$ $\sim\mathcal{N}(0,1)^{d\times p}\,,\qquad s_{i}=\frac{f_{i}U^{\top}}{\sqrt{d}}\in\mathbb{R}^{d}\,$, (F.39)
    and normalize $\tilde{s}_{i}=s_{i}/\lVert s_{i}\rVert_{2}$. Initialize the selected set $S_{0}$ with one random index and iteratively add
    $i_{t+1}=\operatorname*{arg\,max}_{i\in\mathrm{pool}\setminus S_{t}}\,\min_{j\in S_{t}}\lVert\tilde{s}_{i}-\tilde{s}_{j}\rVert_{2}$ (F.40)
    until $|S_{T}|=k$. Return the indices in $S_{T}$.
For each model trained on $N$ samples, performance is measured by the relative OOD error (Eqn. F.19), approximated with the 205 held-out samples.
Fig. 6 compares the model trained on $N$ samples from the optimal distribution with the two active-learning-based models. The results show that the model trained on the optimal distribution performs comparably to the active learning methods. A key advantage of the AMA approach is that it allows sampling beyond the size of the fixed data pool, whereas pool-based active learning methods are restricted to selecting at most the available pool samples. We remark that generally within the field of active learning and adaptive sampling, one cannot expect to always outperform nonadaptive/i.i.d. sampling, especially in high dimensions. The target mapping and choice of deployment regime $\mathbb{Q}$ will determine the gains, if any, of active sampling [2].
Computational resources.
All experiments were conducted on AWS using Galaxy’s Deep Learning AMI GPU, PyTorch 2.0.1, and Ubuntu 20.04-prod-ada4rwvf26mda with an EC2 instance type of g5.12xlarge. The experiments for AMA and the two different active learning methods were run in parallel, with the QbC method taking the longest at 11 hours. Thus, a total of approximately 11 GPU-hours (and corresponding CPU usage) were required for all independent runs with four GPUs running in parallel.